A researcher at a major clinic wishes to estimate the proportion of the adult population of the United States that has a certain health condition. How large of a sample is needed in order to be 95% confident that the sample proportion will not differ from the true proportion by more than 3%?

Group of answer choices

1113

1068

752

1344

1535
Answer #1

Solution :

Given that,

\hat p= 0.5

1 - \hat p = 1 - 0.5 = 0.5

margin of error = E = 3% = 0.03

At 95% confidence level the z is ,

\alpha = 1 - 95% = 1 - 0.95 = 0.05

\alpha / 2 = 0.05 / 2 = 0.025

Z\alpha/2 = Z0.025 = 1.96 ( Using z table )

Sample size = n = (Z\alpha/2 / E)2 * \hat p * (1 - \hat p )

= (1.96 / 0.03)2 * 0.5 * 0.5

= 1068

Sample size =1068